# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v1

@@ Appendix B Proof of Lemma 1
Fix a realized dataset $\mathcal{D}_{n}$ and a node index $v$. All probability statements are taken with respect to the tree-generating randomization. For each tree $\theta$, let $S^{(m)}_{\theta,v}\subset\{1,\dots,p\}$ denote the candidate variable set drawn at node $v$ under candidate-set size $m$.
@@ Appendix C Proof of Theorem 2
Fix a prediction point $x$ and a realized dataset $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. All expectations and covariances are taken with respect to the tree-generating randomization.
Design-based second-moment convention. The quantities $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X_{i})$ are treated as fixed conditional second moments evaluated at the observed covariate values and enter the covariance expansion only as scaling factors.
@@
Step 3: Conclusion. Combining Step 1 and Step 2 yields $C_{T}(x)>0$. The limit statement follows from Theorem 1, since $\lim_{B\to\infty}\operatorname{Var}(\hat{f}_{B}(x))=C_{T}(x)$. ∎
@@ Appendix D Proof of Proposition 2
Fix a prediction point $x$. Consider two forest designs that differ only in the candidate-set size at each split, with $q^{\prime}<q$. Let
$Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta}],\qquad Z_{\theta}^{(q^{\prime})}(x)=\mathbb{E}[T_{\theta}^{(q^{\prime})}(x)\mid\mathcal{I}_{\theta}]$

### v2

@@ Appendix B Proof of Lemma 1
Fix a realized covariate configuration $X$ and a node index $v$. All probability statements are taken with respect to the tree-generating randomization. For each tree $\theta$, let $S^{(m)}_{\theta,v}\subset\{1,\dots,p\}$ denote the candidate variable set drawn at node $v$ under candidate-set size $m$.
@@ Appendix C Proof of Theorem 2
Fix a prediction point $x$ and a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$. All expectations and covariances are taken with respect to the joint distribution of $(Y\mid X)$ and the independent tree-generating randomizations. For this proof, we suppress $X$ throughout for notational clarity; all statements are understood to be conditional on the fixed covariate configuration consistent with the main text.
Conditional second moments. The quantities $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$ denote the conditional outcome variances evaluated at the realized covariate values and enter the covariance expansion as deterministic scale factors once $X$ is fixed.
@@
Step 3: Conclusion. Combining Step 1 and Step 2 yields $C_{T}(x)>0$. The limit statement follows from Theorem 1, since $\lim_{B\to\infty}\operatorname{Var}(\hat{f}_{B}(x)\mid X)=C_{T}(x).$. ∎
@@ Appendix D Proof of Proposition 2
All probability statements are taken with respect to the tree-generating randomization under fixed $X$. Fix a prediction point $x$. Consider two forest designs that differ only in the candidate-set size at each split, with $q^{\prime}<q$. Let
$Z_{\theta}^{(q)}(x)=\mathbb{E}[T_{\theta}^{(q)}(x)\mid\mathcal{I}_{\theta},X],\qquad Z_{\theta}^{(q^{\prime})}(x)=\mathbb{E}[T_{\theta}^{(q^{\prime})}(x)\mid\mathcal{I}_{\theta},X]$